### v1

# DMAligner: Enhancing Image Alignment via Diffusion Model Based View Synthesis

Abstract
Image alignment is a fundamental task in computer vision with broad applications. Existing methods predominantly employ optical flow-based image warping. However, this technique is susceptible to common challenges such as occlusions and illumination variations, leading to degraded alignment visual quality and compromised accuracy in downstream tasks. In this paper, we present DMAligner, a diffusion-based framework for image alignment through alignment-oriented view synthesis. DMAligner is crafted to tackle the challenges in image alignment from a new perspective, employing a generation-based solution that showcases strong capabilities and avoids the problems associated with flow-based image warping. Specifically, we propose a Dynamics-aware Diffusion Training approach for learning conditional image generation, synthesizing a novel view for image alignment. This incorporates a Dynamics-aware Mask Producing (DMP) module to adaptively distinguish dynamic foreground regions from static backgrounds, enabling the diffusion model to more effectively handle challenges that classical methods struggle to solve. Furthermore, we develop the Dynamic Scene Image Alignment (DSIA) dataset using Blender, which includes 1,033 indoor and outdoor scenes with over 30K image pairs tailored for image alignment. Extensive experimental results demonstrate the superiority of the proposed approach on DSIA benchmarks, as well as on a series of widely-used video datasets for qualitative comparisons. Our code is available at https://github.com/boomluo02/DMAligner.
footnotetext: 22footnotemark: 2 Corresponding Author.
1 Introduction
The alignment of consecutive frames is a critical task in computer vision, supporting a wide range of applications such as High Dynamic Range (HDR) imaging [53, 6, 8], video stabilization [26], image deblurring [58], and super-resolution [5]. Precise alignment ensures the accurate merging of information from consecutive frames, leading to enhanced image quality and better performance in subsequent tasks [37, 51, 52]. Despite its significance, the task of image alignment remains challenging due to various factors, including scene dynamics, large camera motion, and varying lighting conditions.
Figure 1: (a) Conventional image alignment based on optical flow and image warping, resulting in ghosting artifacts and occlusion. (b) Our DMAligner directly generate the complete alignment image via diffusion-based view synthesis.
Recent advancements in deep learning have led to substantial progress in neural network-based alignment methods. Generally, image alignment can be divided into three levels: global-level, mid-level, and pixel-level alignment. Global-level alignment typically involves estimating a homography matrix, which can be accomplished using unsupervised deep learning frameworks [23, 25]. This technique effectively aligns images under conditions of planar scenes or when the camera undergoes basic rotational or translational motion. Moreover, through global-to-local homography flow refinement, it can be naturally extended to local mesh-grid homography estimation, allowing it to overcome the limitations of a single homography [23]. However, global- and mid-level methods are restricted in their ability to address more complex scenarios, like parallax, non-rigid deformations, or significant foreground motion, since they assume single or multiple global transformations across the entire image.
Another prominent technique involves the pixel-level approaches with optical flow and image warping [11, 2, 49, 63]. Despite progress in image alignment, current optical flow-based methods face challenges that limit their effectiveness in practical applications. A major issue is the sensitivity of optical flow methods to brightness changes. Variations in lighting conditions, such as shadows, reflections, or changes in ambient light, can significantly distort the brightness constancy assumption, leading to incorrect motion estimation. Besides, the warping process inherent in optical flow-based alignment can introduce various artifacts. Inaccuracies in the motion field can lead to distortions such as stretching, compressing, or tearing of image regions. Additionally, the many-to-one mapping inherent in image warping on occlusion regions inevitably results in ghosting artifacts, as depicted in Fig. 1 (a), even with precise optical flow estimation. Such artifacts not only degrade the visual quality of the aligned frames but also affect the accuracy of subsequent processing tasks.
In this paper, we introduce DMAligner, a diffusion-based framework for image alignment through alignment-oriented view synthesis. As depicted in Fig. 2, the desired image alignment result represents a scenario where the camera pose stays fixed at $P_{1}$ while time advances from $t_{1}$ to $t_{2}$. DMAligner is designed to handle the challenges in image alignment from a new perspective with a generation-based solution (see Fig. 1 (b)), exhibiting strong capabilities and avoiding issues associated with homography and flow-based image warping. Specifically, we first create Dynamic Scene Image Alignment (DSIA) dataset, designed to simulate typical image alignment challenges including camera shifts, moving objects, and fluctuating illumination. Moreover, we propose the Dynamics-aware Diffusion Training approach on the Latent Diffusion Model [38] for learning conditional image generation. The key idea is to enhance the network’s capacity to recognize dynamic information during diffusion model training. This is realized through the Dynamics-aware Mask Producing (DMP) module, which effectively discriminates dynamic foreground regions from static backgrounds within the hidden features across frames. Additionally, we validate the effectiveness of DMAligner through extensive testing on synthetic datasets, including the proposed DSIA and Sintel [3], as well as on the widely-used real-world dataset DAVIS [4]. Both the quantitative and qualitative results demonstrate significant enhancements over contemporary methods. The contributions are summarized as follows:
  • We develop DSIA dataset, the first large-scale dataset specific for image alignment, which simulates the typical challenges including camera motion, moving objects, and changes in illumination.
  • We present DMAligner, a novel diffusion-based framework for image alignment via alignment-oriented view synthesis. The central idea is to enhance the network’s capability to capture dynamic information during diffusion model training, with the assistance of our DMP module.
  • The proposed method achieves state-of-the-art performance on DSIA benchmarks and also demonstrates the superior generalization capability on both synthetic and real-world datasets.
Figure 2: Overview of DSIA dataset generation. The ground truth image is rendered by setting the time to $t_{2}$ and the camera pose to $P_{1}$, where the background statics mirror those in $I_{1}$, and the dynamics are akin to $I_{2}$, with slight visualization variances caused by pose changes. As a result, $I_{gt}$ can be regarded as the alignment-oriented view synthesis with reference from $I_{2}$ to $I_{1}$ (i.e., $I_{2}$$^{\prime}$ for $I_{2}$ to $I_{1}$ alignment).
2 Related Work
Image Alignment.
Traditional approaches primarily rely on feature-based methods, where keypoints are detected and matched across images, followed by geometric transformation estimation to align the images [22, 62]. In the deep learning era, DCNNs [68] have revolutionized image alignment in feature space, which is widely applied in video super-resolution [46], video denoising [61], and burst image restoration [10]. In image space, Global- and middel-level alignments typically involve estimating a homography matrix [25] and local mesh-grid homography [23], respectively. But they cannot handle complex scenarios, involving parallax, non-rigid deformations, or significant foreground motion. A notable advancement in image alignment is the adoption of optical flow-based image warping techniques. Optical flow calculates pixel-wise motion between consecutive images, enabling the warping of one image to align with another. With the advances of recent optical flow approaches [36, 41, 29, 7, 33], this method has found widespread use across a variety of video-based applications [27, 1, 2, 63, 11, 55]. Despite their success, warping in occlusion regions inevitably leads to ghosting artifacts, even with accurate flow, as shown in Fig. 1. In contrast, our DMAligner addresses the challenges in image alignment from a new perspective, performing alignment-oriented view synthesis using a diffusion-based framework.
Figure 3: Overview of our DMAligner. Instead of using the discriminative learning paradigm for optical flow estimation and image warping, our framework employs a generative approach to achieve image alignment with a diffusion model. Dynamics-aware Mask Producing (DMP) module is crucial for providing dynamic information, essential for performing the Dynamics-aware Diffusion Training process in this task.
Diffusion Model.
Diffusion models own prominent generative performance for synthesis [56, 14], and various approaches have been developed to improve the performance or efficiency, such as DDPM [14], DDIM [42], LDM [38] and DDNM [50] etc. Different applications have shown good performances when empowered with diffusion models, including but not limited to, optical flow estimation [39, 28], shadow removal [30], image rectangling [67], homography estimation [20], rolling shutter removal [57], low-light image enhancement [15, 16] and camouflaged object detection [65]. In this work, we propose an end-to-end diffusion-based framework that specifically handles cross-frame dynamic information for image alignment.
3 Method
3.1 DSIA Dataset Generation
[…]
Specifically, inspired by PointOdyssey [66] and Kubric [13], we assemble 25 characters and 100 objects with diverse shapes and materials from several open-source sources, including BlenderKit, Mixamo, GSO [9], and PartNet [35]. These characters are animated using real-world motion capture data [34], while the locations and rotations of object poses are randomly adjusted to simulate realistic motion [13]. Camera motion is controlled along predefined trajectories extracted from [21]. By collecting and animating these assets (characters, objects, and cameras), we generate 3D virtual scenes that closely mimic real-world dynamics. These scenes are then rendered with the Blender engine, producing image sequences $\{I_{1},I_{2},\cdots,I_{N}\}$, in which characters, objects, and the camera all move simultaneously. To maximize the diversity of the generated images, we randomly introduce HDR environments with various materials, textures, and lighting conditions, similar to those in [48].
We systematically rearrange the sequencing of the camera trajectories relative to character and object motions, capturing dynamic scenes from various positions along the trajectories, and rendering to generate virtual viewpoint images tailored for image alignment. Considering two consecutive images, $I_{1}$ and $I_{2}$ involve both global camera motion $\mathcal{H}$ and local foreground movement $\mathcal{F}$. Their transformations can be described using the traditional warp operation $\mathcal{W}$:
$I_{1}=\mathcal{W}(I_{2})+\mathcal{T}=\mathcal{H}(\mathcal{F}(I_{2}))+\mathcal{T},$ (1)
where $\mathcal{T}$ represents the changes in image texture. In our approach, we capture the dynamic scene at the time of $I_{2}$ from the camera viewpoint of $I_{1}$, generating the aligned image $I_{gt}$ as Ground Truth. Compared to $I_{1}$, $I_{gt}$ involves local foreground movement $\mathcal{F}$, while compared to $I_{2}$, $I_{gt}$ involves global camera motion $\mathcal{H}$:
$I_{gt}=\mathcal{H}(I_{2})=\mathcal{F}(I_{1})+\mathcal{T}.$ (2)
The dataset is divided into four subsets based on varying camera and foreground movements.
3.2 Diffusion Model for Image Alignment
3.2.1 Overview.
The training and inference pipelines of our DMAligner are depicted in Fig. 3. The entire process is formulated in the latent space using LDM[38]. The core insight is to equip the network with the ability to capture dynamic information during diffusion model training. This is accomplished by the Dynamics-aware Mask Producing (DMP) module, which effectively differentiates dynamic foreground regions from static backgrounds within the cross-frame hidden features. We elaborate on the model training and inference details in the following subsections.
3.2.2 Dynamics-aware Diffusion Training.
Given a set of image data $I_{1}$, $I_{2}$, and $I_{gt}$, we follow LDM [38] to transform these high-resolution images into the latent representational space using a perceptual compression model based on the encoder-decoder architecture, and then reconstruct them back into RGB images. Specifically, for $I_{gt}$, the encoder $\mathcal{E}$ encodes it into the latent representation $V_{gt}$, and the decoder $\mathcal{D}$ reconstructs it into the RGB image:
$\tilde{I}_{gt}=\mathcal{D}(V_{gt})=\mathcal{D}(\mathcal{E}(I_{gt})).$ (3)
The same process is applied to $I_{1}$ and $I_{2}$, resulting in latent representations $V_{1}$ and $V_{2}$. We use $V_{1}$ and $V_{2}$ as the condition in the latent representational space and train the diffusion model $\theta$ to learn sampling from standard Gaussian noise, generating the latent that matches the distribution of $V_{gt}$. Our training scheme consists of four components: Forward Diffusion, Conditional Learning, Dynamics-aware Mask Producing, and Loss Function.
Forward Diffusion: As illustrated in Fig. 3, we apply different timesteps $t$ to add varying levels of noise to $V_{gt}$ (denoted as $x_{0}$), generating a series of noise representations $x_{t}\in\{x_{1},x_{2},\cdots,x_{T}\}$. $x_{t}$ indicates any noise representation at a specific timestep $t$, and as $t$ increases, the level of added noise grows, leading to an increase in the entropy of the distribution of $x_{t}$. Ultimately, the distribution becomes standard Gaussian noise, $x_{T}\sim\mathcal{N}(0,1)$, defined as:
$x_{t}\in\{x_{1},x_{2},\cdots,x_{T}\},\ x_{T}\sim\mathcal{N}(0,1),\ x_{0}=V_{gt}.$ (4)
The forward diffusion process typically follows the approach of [14], where Gaussian noise is incrementally added to the data sample based on a Markov chain. This allows for the diffusion from $x_{0}$ to any timestep $x_{t}$ in one step using the following equation:
$q(x_{t}\mid x_{0})=\mathcal{N}(x_{t}\mid\sqrt{\bar{\alpha}_{t}}x_{0},(1-\bar{\alpha}_{t})I),$ (5)
where $\bar{\alpha}_{t}$ represents the cumulative product of a set of pre-defined noise variance schedules $\{\beta_{1},\beta_{2},\cdots,\beta_{t}\}$, which ensures that the noise incrementally increases with $t$, defined as $\bar{\alpha}_{t}:={\textstyle\prod_{i=1}^{t}}\alpha_{i}={\textstyle\prod_{i=1}^{t}}(1-\beta_{i})$.
Dynamics-aware Mask Producing: Notably, our approach explicitly leverages the motion foreground mask to guide the network $\theta$ in learning denoising process. This guidance makes the network focus on the moving foreground objects, enabling the better reconstruction of aligned images with enhanced local details. The model relies on our DMP to generate motion foreground masks during training. Specifically, we first compute the correlation $\mathcal{C}(V_{1},V_{2})$ between $V_{1}$ and $V_{2}$ to extract sufficient motion information, then input $\mathcal{C}(V_{1},V_{2})$ into the mask predictor $\mathcal{M}$ to derive the predicted mask $\tilde{M}_{pred}$, as shown in Eq. 6:
$\mathcal{C}(V_{1},V_{2})$ $=\sum_{o\in\Theta}\left\langle V_{1}(u),V_{2}(u+o)\right\rangle,$ (6)
$\tilde{M}_{pred}$ $=\mathcal{M}(\mathcal{C}(V_{1},V_{2})),$
where $u$ represents the spatial coordinates of $V_{1}$, $\Theta$ denotes the neighborhood of coordinate $u$, and $o$ is the coordinate offset. Importantly, when using the foreground mask to guide the denoising process, we also aim to ensure that network notices the transition regions between foreground and background. To achieve smooth transitions at the edges of foreground objects, we apply the dilation operation $d_{r}$ to expand the produced mask $r$ times.
$V_{M}=V_{2}\times d_{r}(\tilde{M}_{pred})+V_{1}\times\{1-d_{r}(\tilde{M}_{pred})\},$ (7)
as shown in Eq. 7, the expanded foreground mask $d_{r}(\tilde{M}_{pred})$ is used to fuse the background of $V_{1}$ with the foreground of $V_{2}$, creating the mixed latent representation $V_{M}$, which serves as a condition to guide the diffusion network in generating the aligned image output.
Conditional Learning: As shown in Fig. 3, we concatenate $V_{1}$, $V_{2}$ and $V_{M}$ along the channel dimension to form the condition $x_{cond}$, then randomly select $x_{t}$ from the noise representations $\{x_{1},x_{2},\cdots,x_{T}\}$, allowing the diffusion model $\theta$ to learn the denoising process. This enables $\theta$ to handle any $x_{t}$ at an arbitrary timestep $t$, producing a denoised output $\hat{x}_{0}$. Similar to approaches in [45, 18, 28], we use the “predict $x_{0}$” denoising setting in our model. The condition $x_{cond}$ is concatenated with $x_{t}$ and the encoded timestep $t$ through positional embedding (PE) [47], and the result input is sent to the diffusion model to obtain the denoised result $\tilde{x}_{0}$:
$\tilde{x}_{0}=\theta([x_{t},x_{cond},{\rm PE}(t)]),$ (8)
then $\tilde{x}_{0}$ is reconstructed to become $\tilde{I}_{pred}$ by decoder $\mathcal{D}$, which can be expressed as $\tilde{I}_{pred}=\mathcal{D}(\tilde{x}_{0})$. Experiments demonstrate that “predict $x_{0}$” is more effective for image alignment compared to others.
Figure 4: Qualitative comparisons between DPFlow [36], COGS [17], GenWarp [40] and our DMAligner on our DSIA dataset.
Table 1: Comparison of network inputs among DPFlow [36], COGS [17], GenWarp [40], and our DMAligner. Other optical flow networks (PWCNet [43], RAFT [44], FlowFormer++ [41], FlowDiffuser [28]) have the same input structure as DPFlow.
Method Main Input Structure
DPFlow [36] 2 images ($I_{1},I_{2}$)
COGS [17] 2 images ($I_{1},I_{2}$) +
Depth ($D_{1},D_{2}$) + Mask ($M_{1},M_{2}$)
GenWarp [40] Single image ($I_{2}$) + Camera Intrinsics ($K$) +
Depth ($D_{2}$) + Camera Motion ($P_{I_{2}\to I_{1}}$)
DMAligner 2 images ($I_{1},I_{2}$)
Table 2: Quantitative comparison on our DSIA dataset across four subsets (LcLf, LcSf, ScLf, and ScSf) using PSNR and SSIM metrics, higher scores indicate better performance. The average (“Avg”) across all subsets summarizes overall results. indicates that occlusion and ghost regions are excluded during metric computation, as determined by forward-backward consistency check [54].
Methods LcLf LcSf ScLf ScSf Avg.
[…]
Methods based on Flow Warping
PWCNet [43] 16.53 0.74 16.40 0.73 20.53 0.74 21.01 0.76 18.62 0.74
RAFT [44] 18.22 0.74 18.34 0.74 22.46 0.77 22.95 0.76 20.49 0.75
FlowFormer++ [41] 21.02 0.75 20.97 0.76 22.84 0.77 23.94 0.79 22.19 0.77
FlowDiffuser [28] 20.18 0.75 20.33 0.75 23.39 0.78 23.91 0.79 21.95 0.77
DPFlow [36] 20.78 0.76 20.50 0.74 23.94 0.78 24.61 0.80 22.46 0.77
PWCNet [43] 22.03 0.76 21.15 0.76 24.63 0.79 24.74 0.77 23.14 0.77
RAFT [44] 22.99 0.78 23.15 0.77 25.07 0.80 25.09 0.79 24.08 0.78
FlowFormer++ [41] 24.19 0.79 24.24 0.78 25.61 0.81 25.74 0.81 24.95 0.80
FlowDiffuser [28] 23.82 0.78 23.81 0.77 25.78 0.81 25.89 0.82 24.82 0.80
DPFlow [36] 24.30 0.79 24.84 0.79 25.99 0.82 26.04 0.82 25.29 0.81
Methods based on View Synthesis
COGS [17] 16.35 0.66 16.84 0.69 17.01 0.69 18.21 0.71 17.10 0.69
SD-Inpaint [59] 22.14 0.75 22.92 0.76 25.42 0.79 25.77 0.81 24.06 0.78
GenWarp [40] 22.62 0.76 23.15 0.76 25.32 0.78 25.84 0.81 24.23 0.78
DMAligner(Ours) 24.91 0.80 26.90 0.81 27.25 0.81 27.64 0.83 26.67 0.81
Figure 5: Qualitative comparisons between DPFlow [36], GenWarp [40] and our DMAligner on Sintel dataset.
Figure 6: Qualitative comparisons between DPFlow [36], GenWarp [40] and our DMAligner on the real-world DAVIS dataset.
Figure 7: Comparison of HDR results using flow-based method HDRFlow [53], DPFlow [36] and our DMAligner. Our DMAligner significantly reduces ghosting and background distortions, especially around occluded regions and moving objects.
Table 3: Quantitative comparison on Sintel and DAVIS dataset. Due to the lack of ground truth for precise evaluation, we compute LPIPS [64] and DreamSim [12] between $I_{pred}$ and $I_{1}$. Lower scores indicate better performance.
Methods Sintel (Clean) Sintel (Final) DAVIS Avg.
$downarrow LPIPS\downarrow DreamSim\downarrow$ $downarrow LPIPS\downarrow DreamSim\downarrow$
Methods based on Flow Warping
PWCNet [43] 0.251 0.146 0.177 0.109 0.217 0.102 0.215 0.119
RAFT [44] 0.245 0.141 0.171 0.102 0.219 0.104 0.212 0.116
FlowFormer++ [41] 0.264 0.150 0.184 0.113 0.224 0.106 0.224 0.123
FlowDiffuser [28] 0.258 0.141 0.179 0.103 0.211 0.090 0.216 0.111
DPFlow [36] 0.255 0.138 0.178 0.101 0.209 0.087 0.214 0.109
Methods based on View Synthesis
COGS [17] 0.411 0.172 0.378 0.131 0.398 0.121 0.396 0.142
SD-Inpaint [59] 0.269 0.164 0.271 0.142 0.241 0.114 0.260 0.140
GenWarp [40] 0.264 0.152 0.251 0.130 0.234 0.108 0.250 0.130
DMAligner(Ours) 0.246 0.123 0.186 0.117 0.202 0.084 0.211 0.108
Loss Function: Our model directly learns to denoise and reconstruct a latent representation $\tilde{x}_{0}$ from $x_{t}$ that matches the distribution of $V_{gt}$. Then, it is decoded by $\mathcal{D}$ into the RGB image $\tilde{I}_{pred}$, supervised by $I_{gt}$ in the image domain to ensure reliable and accurate output. Additionally, the trainer learns the motion foreground mask $\tilde{M}_{pred}$, which is then upsampled to obtain $\hat{M}_{pred}$, helping the model to focus on regions with significant variance. The denoising loss function, namely Denoising Loss, is defined as follows:
$\hat{M}_{pred}=Upsample\{d_{r}(\tilde{M}_{pred})\},$ (9)
$L_{\rm{Denoising}}$ $=(1-\gamma)\left\|(1-\hat{M}_{pred})(I_{gt}-I_{pred})\right\|_{2}$ (10)
$+\gamma\left\|(\hat{M}_{pred})(I_{gt}-I_{pred})\right\|_{2},$
where $\gamma$ is a variable weighting coefficient, and $d_{r}$ represents the dilation operation applied to the mask $r$ times. We also supervise the mask $\tilde{M}_{pred}$ using the cross-entropy loss, called as Mask Loss:
$L_{\rm{Mask}}=-M_{gt}\log(\hat{M}_{pred})-(1-M_{gt})\log(1-\hat{M}_{pred}).$ (11)
The total loss is:
$L_{\rm{Total}}=\lambda_{1}L_{\rm{Denoising}}+\lambda_{2}L_{\rm{Mask}}.$ (12)
In our experiments, the parameters $r$ (in $d_{r}$), $\lambda_{1}$, and $\lambda_{2}$ are set to 2, 2, and 0.1, respectively.
3.2.3 Dynamics-aware Denoised Inference.
After training, the diffusion model $\theta$ performs inference by denoising a standard Gaussian noise $x_{T}\in\mathcal{N}(0,1)$, conditioned on $V_{1}$, $V_{2}$ and $V_{M}$. This process reconstructs a reliable and accurate alignment result through recurrent denoising. Specifically, the model $\theta$ takes $x_{T}$, the condition, and the timestep $T$ as inputs according to Eq. 8, producing an initial coarse denoised output $\hat{x}_{0}^{(T)}$. Then, following the non-Markovian forward process [42] parametrized by $\sigma$ as shown in Eq. 13, it generates $x_{T-1}$ at timestep T1T-1. This process is repeated recurrently, producing a sequence of progressively refined denoised results $\hat{x}_{0}^{(T)},\hat{x}_{0}^{(T-1)},\cdots,\hat{x}_{0}^{(1)}$. Finally, $\hat{x}_{0}^{(1)}$ is passed through the decoder $\mathcal{D}$ to output the aligned image $I_{pred}$. The non-Markovian forward process is defined as:
$x_{t-1}=\sqrt{\alpha_{t-1}}\hat{x}_{0}^{(t)}+\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}\hat{\epsilon}_{t}+\sigma_{t}x_{T},$ (13)
where $\hat{x}_{0}^{(t)}$ is the prediction of $x_{0}$ from $x_{t}$ using the diffusion model $\theta$, and $\hat{\epsilon}_{t}$ is the approximate noise calculated by
$\hat{\epsilon}_{t}=\frac{x_{t}-\sqrt{\alpha_{t}}\hat{x}_{0}^{(t)}}{\sqrt{1-\alpha_{t}}}.$ (14)
In the recurrent update process described above, the step size $\Delta$ is not fixed and can be adjusted to other values. By setting an appropriate step size $\Delta$, inference can be accelerated while still allowing incremental refinement of the results. The iterative update process can be expressed as: $\{x_{T}\to\hat{x}_{0}^{(T)}\}\to\{x_{T-\Delta}\to\hat{x}_{0}^{(T-\Delta)}\}\to\cdots\to\{x_{1}\to\hat{x}_{0}^{(1)}\}$. In our empirical analysis, we set $T=1000$ and $\Delta=20$.
4 Experiments
4.1 Input Structure
As shown in Tab. 1, the input requirements for DPFlow [36], COGS [17], GenWarp [40], and our DMAligner are explained. Both DPFlow and DMAligner only require two consecutive image frames $I_{1}$ and $I_{2}$ for optical flow estimation and view synthesis, respectively. On the other hand, COGS relies on Gaussian splatting and requires not only the two input images $I_{1}$ and $I_{2}$, but also foreground masks ($M_{1},M_{2}$) and depth maps ($D_{1},D_{2}$), with the masks predicted by FC-CLIP [60] and depth estimated by Marigold [19]. These additional inputs are crucial for its 3D scene reconstruction. Although GenWarp is described as using only a single image $I_{2}$ as input, it actually relies on both $I_{1}$ and $I_{2}$ to obtain the relative camera motion $P_{I_{2}\to I_{1}}$, which can be derived from the optical flow $F_{I_{2}\to I_{1}}$ estimated by DPFlow, filtering out local motion to capture the overall camera movement [31, 32, 24]. Besides, it also needs camera intrinsics ($K$), depth ($D_{2}$) from Marigold. Therefore, GenWarp also relies on information from both $I_{1}$ and $I_{2}$. Unlike GenWarp and COGS, our DMAligner does not require additional geometric information such as camera intrinsics, depth maps, or pose transformations, making it a simpler and more direct approach for alignment.
4.2 Implementation Details
[…]
4.3 Comparison with Existing Methods
4.3.1 Results on DSIA.
Tab. 2 compares our DMAligner with some flow-warping methods and view-synthesis methods. In our DSIA dataset, all flow-warping methods are trained using rendered ground truth optical flow. We then align $I_{2}$ to $I_{gt}$ using back warping and compute PSNR and SSIM between $I_{pred}$ and $I_{gt}$. To address occlusion and ghosting artifacts from warping, these regions (determined by forward-backward consistency check [54]) are excluded from the metrics calculation, as marked by . Fig. 4 shows qualitative visual results on the DSIA test set, our DMAligner most closely matches the ground truth.
4.3.2 Results on Sintel and DAVIS.
We conduct qualitative comparisons on and the synthetic Sintel [3] dataset and the realistic DAVIS [4] dataset to evaluate the cross-domain performance of DMAligner and DSIA dataset. All competing methods and our DMAligner are trained on DSIA, and alignment results are directly tested on the Sintel Clean, Sintel Final training sets, and the DAVIS dataset. As there is no ground truth for precise evaluation, we calculate the feature distances between $I_{pred}$ and $I_{1}$ using LPIPS [64] and DreamSim [12]. Tab. 3 presents the quantitative results on Sintel and DAVIS, where DMAligner achieves the best average LPIPS and DreamSim scores. Fig. 6 shows examples from the DAVIS dataset, and Fig. 5 displays results from the Sintel test set. These results demonstrate that DMAligner delivers exceptional image alignment performance in both synthetic and real-world scenarios.
4.3.3 Results on Downstream HDR Task.
In HDR imaging [53], LDR frames with alternating exposures (e.g., $L_{t-1}$$L_{t}$$L_{t+1}$) are aligned using flow warping before fusion. We replace this step with DMAligner, which aligns foreground and background more robustly, generating warped frames ($L_{t+1\to t}$, $L_{t-1\to t}$). Importantly, the fusion network and overall pipeline remain unchanged—only the aligned inputs differ. In Fig. 7, our method produces fewer artifacts, demonstrating strong generalization with existing HDR fusion network, which is expected to improve other downstream tasks as well.
Figure 8: Ablation experiments of DMP. The results of DMAligner without or with DMP are given, as well as the mask prediction $\hat{M}_{pred}$ and mask labels $M_{gt}$.
4.4 Ablation Study
In Tab. 4, we explore the impact of predict methods and DMP. Comparing (a)&(b) and (c)&(d)&(e), “predict $x_{0}$” is more effective for image alignment, as it directly targets image reconstruction. comparing (b)&(e), DMP resultes in more accurate alignment, as it allows the model to better focus on relevant areas. Fig. 8 also demonstrates that DMP enables the network to focus on the local dynamic foreground areas, improving the foreground details of the generated aligned images and achieving superior alignment.
Table 4: Ablation studies of predict methods and our DMP.
Exp. Predict DMP Avg
Methods PSNR SSIM
(a) pred $\hat{\epsilon}_{t}$ w/o 24.82 0.77
(b) pred $x_{0}$ w/o 25.12 0.79
(c) rectified flow w/ 25.98 0.80
(d) v-prediction w/ 26.16 0.81
(e) pred $x_{0}$ w/ 26.67 0.81
5 Conclusion
This work represents a significant paradigm shift in image alignment by introducing a generation-based solution. The proposed DMAligner framework, a diffusion-based approach for image alignment through alignment-oriented view synthesis, effectively addresses challenges such as ghosting artifacts, occlusions, and illumination changes that previous methods struggle with. Notably, a Dynamics-aware Mask Producing (DMP) module is introduced to enable the diffusion model to better handle dynamic foreground regions that are prone to errors. Additionally, the Dynamic Scene Image Alignment (DSIA) dataset is developed specifically for training. Extensive experiments demonstrate that DMAligner outperforms previous methods with unmatched performance and exhibits strong cross-domain capabilities, excelling in both synthetic and real-world scenarios.